The second of three numbers is 1 less than the first. The third number is 5 less than twice the second. If the third number exceeds the first number by 12, find the three numbers

Answer by lwsshak3(11628)

The second of three numbers is 1 less than the first. The third number is 5 less than twice the second. If the third number exceeds the first number by 12, find the three numbers.
**
let x=1st number
x-1=2nd number
2x-5=3rd number
..
2x-5-x=17
x=17
x-1=16
2x-5=27
1st number=17
2nd number=16
3rd number=27